# Unharmful Backdoor-based Client-side Watermarking in Federated Learning

### v1

@@ 5.7 Evaluation with Different Variants
We further investigate the impact of different variants (e.g., hyperparameters and configuration parameters) on the performance of Sanitizer, as well as its sensitivity to the changes in them. Our ablation studies mainly examine the effectiveness and efficiency of Sanitizer under two key factors in our setting: (i) varying numbers of reverse engineering epochs per round and (ii) different backdoor subnet sizes. For each factor, we only vary this single factor while keeping all other setups consistent with the default setting of the evaluation in Section 5.2. All the ablation experiments in this subsection are performed on CIFAR10 using ResNet18.
@@
Impact of Different Round-spread Reverse Engineering Epochs on Efficiency. As shown in Figure 9, we investigate the metrics about efficiency across different round-spread reverse engineering epochs per round: (i) the GPU_Mem (MB) consumed for processing a single client-submitted model during server-side operations, and (ii) the average Time (seconds) required for server-side operations per round. We also report the baseline values under the same settings for reference. These values are fixed, as the baseline does not involve the two parameters. According to the results, GPU memory utilization remains consistent, ranging from approximately 190 MB to 194 MB across all epoch settings, which is notably lower than the baseline. This stability suggests that GPU memory usage of Sanitizer is insensitive to the number of reverse engineering epochs per round. On the other hand, the time required for server-side operations per round shows a clear upward trend, escalating from 19.44 seconds at 1 epoch to 28.12 seconds at 2 epochs, and further rising to 53.73 seconds at 4 epochs. Such a trajectory in the figures stems from the inherently positive relationship between time and the number of reverse engineering epochs per round. Obviously, it persistently remains below the baseline value under varying epochs. The results show that reducing the number of round-spread reverse engineering epochs can save the server-side operation time, further validating Sanitizer’s optimization for efficiency.
@@
Impact of Different Round-spread Reverse Engineering Epochs on Effectiveness. As illustrated in Figure 10, we investigate the WDR and the ASR under different reverse engineering epochs per round. Specifically, both the baseline and Sanitizer consistently maintain a robust and near-perfect WDR over the set of epochs {1, 2, 4}. The ASR follows a similar trend, showing a slight and routine difference between the two approaches and remaining stable as the number of epochs varies. Results in this experiment confirm that different round-spread reverse engineering epochs have virtually no impact on the Sanitizer, which remains effective for any value of this factor. This shows that the effectiveness of Sanitizer is not sensitive to the number of round-spread reverse engineering epochs. In light of the aforementioned analysis pertaining to efficiency and effectiveness, it can be concluded that, even though Sanitizer achieves comparable effectiveness to the baseline varying epochs, the latter is associated with significantly higher costs.
@@
Impact of Different Backdoor Subnet Sizes on Efficiency. Figure 11 reports the results. Our primary objective is to investigate the relationship of efficiency between the backdoor subnet size and the key efficiency metrics. Sanitizer exhibits an increasing trend in both GPU memory utilization and required process time as the backdoor subnet size grows. In particularly, the GPU memory rises from 117 MB to 379 MB, and the process time per round extends from 16.80 to 38.90 seconds. All of these results largely outperform the baseline, demonstrating the intention behind the design of Sanitizer, which aims to enhance efficiency by reducing the size of the extracted backdoor subnet.
@@
Impact of Different Backdoor Subnet Sizes on Effectiveness. Similarly, we systematically vary the backdoor subnet size across three configurations: 10%, 20%, and 40%, and investigate the impact of them on the effectiveness of Sanitizer, keeping other variables constant to ensure fairness in comparisons. Figure 12 presents the WDR results as the backdoor subnet size increases. Sanitizer achieves a near-perfect WDR of 98.77% at the smallest backdoor subnet size (10%) and maintains a high WDR across all configurations of larger subnet sizes (20% and 40%). As for the ASR, Sanitizer shows a much lower ASR than the baseline and a stable trend across three subnet sizes. Hence, while markedly reducing resource and time consumption, Sanitizer achieves comparable performance to the baseline across different backdoor subnet sizes, demonstrating its effectiveness in eliminating harmful backdoor knowledge and enabling contribution demonstration via our unharmful watermarks. In summary, the effectiveness of Sanitizer is not sensitive to the reduction in subnet size brought about by the need to improve efficiency during the FL process.
@@ 6 Conclusions
With the goal of strengthening trustworthy federated learning, we have introduced Sanitizer, a novel and efficient sanitization approach designed to address the harmful effects and conflict issues posed by backdoor-based client-side watermarks for contribution demonstration in practical FL applications. Sanitizer leverages backdoor subnets and round-spread reverse engineering to efficiently reverse the backdoor triggers, then confines them to their own unharmful, client-specific environment before deployment and eventually makes the watermarks function in unharmful ways to successfully verify the contributions. Empirical experiments have demonstrated the effectiveness, harmlessness, and especially the efficiency of Sanitizer. We advocate for further exploration in this direction and believe that Sanitizer will inspire more advanced work for better IP protection in various real-world applications.
@@ A.1 Data Availability Statements
We fully support the principles of open science, aiming to promote transparency, reproducibility, and collaborative research. The datasets (Fashion-MNIST (FMNIST) [64], CIFAR10 [65], and TinyImageNet [66]) used in this study are publicly available and widely recognized as the benchmark in the image classification tasks. The details of these datasets are as follows:
@@
Input: Total communication rounds of FL $T$; Number of clients $K$; Defense data $(x_{d},y_{d})\in\mathcal{D}_{d}$; Unharmful data $x_{u}\in\mathcal{D}_{u}$; Poisoning rate of backdoor embedding $\rho\%$;
Output: Unharmful Watermarked FL-trained Model $\mathcal{F}(\theta_{*})$.
@@
2:  for each round $t=0,1,\dots,T-1$ do
@@
9:        Conduct full Reverse Engineering on $f(\theta_{k})$ for trigger pattern $\Delta_{k}$ recovery and target class $y^{k}_{t}$ detection via Equation 4.  $\triangleright$ Trigger Recovery.
10:     end for
11:     Aggregation.
12:  end for
13:  Server conducts $\max\mathbb{E}_{(\boldsymbol{x}_{d},y_{d})\in\mathcal{D}_{d}}\mathcal{L}(f\left(x_{d}+\Delta;\theta\right),y_{t})$ on FL-trained Model $\mathcal{F}(\theta)$.      $\triangleright$ Unlearning.
14:  Server conducts $\min\mathbb{E}_{(\boldsymbol{x}_{d},y_{d})\in\mathcal{D}_{d}}\mathcal{L}(f\left(x_{d}+\Delta;\theta\right),y_{d})$ and $\min\mathbb{E}_{x_{u}\in\mathcal{D}_{u}}\mathcal{L}(f\left(x_{u}+\Delta;\theta\right),y_{t})$ on FL-trained Model $\mathcal{F}(\theta)$.           $\triangleright$ Relearning.
15:  return  Unharmful Watermarked FL-trained Model $\mathcal{F}(\theta_{*})$.
Input: Total communication rounds of FL $T$; Number of clients $K$; Defense data $(x_{d},y_{d})\in\mathcal{D}_{d}$; Unharmful data $x_{u}\in\mathcal{D}_{u}$; Size of backdoor subnet $s\%$; Epoch of round-spread reverse engineering per round $e$; Poisoning rate of backdoor embedding $\rho\%$;
Output: Unharmful Watermarked FL-trained Model $\mathcal{F}(\theta_{*})$.
@@
2:  for each round $t=0,1,\dots,T-1$ do
@@
11:        $\triangleright$ Conduct round-spread reverse engineering for $e$ epochs on $f_{s}(\theta_{k}^{*})$ for trigger pattern $\Delta_{k}$ and target class $y^{k}_{t}$ and pruning for aggregation.
@@
15:  Server conducts Unharmful Relearning on FL-trained Model $\mathcal{F}(\theta)$ via Equation 7.
16:  return  Unharmful Watermarked FL-trained Model $\mathcal{F}(\theta_{*})$.
  • FashionMNIST: Fashion-MNIST contains 70,000 28x28 grayscale images across 10 fashion categories. The dataset is split into 60,000 training images and 10,000 test images. FashionMNIST is designed as a modern alternative to the classic MNIST dataset. It can be accessed from https://github.com/zalandoresearch/fashion-mnist. It is supported natively by PyTorch and will be downloaded automatically.
  • CIFAR-10: CIFAR10 consists of 60,000 32x32 color images in 10 different classes. The dataset is divided into 50,000 training images and 10,000 test images, and is widely used for evaluating machine learning models on small-scale object recognition tasks. It is publicly available at http://www.cs.toronto.edu/~kriz/cifar.html. It is supported natively by PyTorch and will be downloaded automatically.
  • TinyImageNet: TinyImageNet includes 100,000 64x64 color images across 200 classes. Each class contains 500 training images and 50 validation images. TinyImageNet is a smaller, more manageable subset of the full ImageNet dataset, commonly used for evaluating deep learning models. It can be downloaded from http://cs231n.stanford.edu/tiny-imagenet-200.zip.
@@
communication rounds 100
local training epochs 5
@@
Unharmful learning rate 0.005
@@
weight for unharmful objective 0.2
Table V: Default Parameter Setting in FL with Sanitizer.
@@ A.3 Default Parameter Setting
Table V summarizes the default parameter setting in our empirical evaluation.
@@ A.4 More Details on the Extracted Backdoor Subnet
Based on the default settings, we present a further analysis of the extracted backdoor subnet in two different architectures when applied to CIFAR10, which has a resolution of 32×32 pixels and three color channels (RGB). As shown in Table VI, we compared the number of parameters and FLOPs between the original network and the extracted backdoor subnet. The results clearly demonstrate a significant reduction in computational complexity. This decrease in model size and operations highlights the efficiency gained through subnet extraction, making it more suitable for resource-constrained environments without sacrificing key performance metrics.
@@ A.5 Visual Examples of Trigger Inversion on the FL-trained Model without Sanitizer
We observe that recovering the trigger from an FL-trained model without employing Sanitizer appears to be not feasible. In this scenario, no defense mechanisms are applied during the training phase of the existing backdoor-based client-side watermarked FL, and reverse engineering is only performed on the final FL-trained model after the FL process has completed, prior to deployment. Table VII illustrates that the shortcut, or trigger, associated with each class cannot be successfully recovered in the final FL-trained model. We conjecture that this outcome may be attributed to the aggregation process, which potentially dilutes the backdoor effect, thereby obscuring the true trigger. This dilution likely results from the averaging of model updates across clients, which disperses the influence of individual backdoor triggers and diminishes their impact on the global model. Consequently, the aggregated model fails to retain the malicious characteristics necessary for trigger identification. The results indicate that it seems not possible to recover the specific triggers associated with each target label, highlighting the difficulty of accurately identifying and extracting the embedded backdoors under this condition.
Table VI: Complexity of the extracted backdoor subnet in two different architectures when applied to CIFAR10.
@@
Table VII: The Original Trigger corresponding to each class and its recovered version.
@@
A.6 More Analysis and Discussion
@@
$\bullet$ Why assume a Badnet-style backdoor? We argue that in a scenario where FL clients co-own the jointly trained model, clients tend to embed watermarks in the simplest and most direct manner to demonstrate their contribution, obviating the need for complex and impractical techniques. As clients cannot easily break into the source code and manipulate the standard local learning, the easiest way for them to watermark the model is to conduct data poisoning backdoors. This means they don’t need to expend significant effort (or resources) to demonstrate their contribution; instead, a more practical and straightforward approach for them is implementing a Badnet-like patch replacement that applies a trigger pattern onto an image and changes its label. It is reasonable to assume that they prefer to demonstrate their contribution with minimal cost rather than utilizing more sophisticated backdoor embedding methods. Moreover, we specifically focus on Badnet-like backdoor-based watermarks, as our approach is not designed to defend against backdoors in general, but rather to serve as a safeguard when such backdoors are used as watermarks under specific conditions.
A.7 Limitation and Future Work
Sanitizer has confirmed the promising results of making the backdoor-based client-side watermarks unharmful to enhance the contribution demonstration in FL, but it also comes with several limitations: (a) Sanitizer has not yet optimized the computational cost associated with subsequent step after FL, although it is relatively negligible compared to the whole FL period; (b) the feasibility of applying Sanitizer to Transformer-based architectures, such as ViT, remains unknown for now.
In the future, we will address these challenges and further extend the applicability of Sanitizer. Sanitizer represents the first research effort that directly targets the harmlessness of backdoor-based watermarks in FL. This work not only advances the state-of-the-art for image classification tasks but also paves the way for future exploration across a broader range of applications, including object detection and beyond. We anticipate that the principles behind Sanitizer can be effectively extended to other FL scenarios, driving further innovations in enhancing the IP protection and contribution demonstration of the FL-trained model.

### v2

@@ 5.9 Parameter Studies
We further investigate the impact of different parameter settings on the performance of Sanitizer, as well as its sensitivity to the variation in them. Our parameter studies mainly examine the effectiveness and efficiency of Sanitizer under two key factors in our setting: (i) varying numbers of reverse engineering epochs per round and (ii) different backdoor subnet sizes. For each factor, we only vary this single factor while keeping all other setups consistent with the default setting of the evaluation in Section 5.2. All the parameter experiments in this subsection are performed on CIFAR10 using ResNet18.
@@
Impact of Different Round-spread Reverse Engineering Epochs on Efficiency. As shown in Figure 9, we investigate the metrics about efficiency across different round-spread reverse engineering epochs per round: (i) the GPU_Mem (MB) consumed for processing a single client-submitted model during server-side operations, and (ii) the average Time (s) required for server-side operations per round. We also report the baseline values under the same settings for reference. These values are fixed, as the baseline does not involve the two parameters. According to the results, GPU memory utilization remains consistent, ranging from approximately 190 MB to 194 MB across all epoch settings, which is notably lower than the baseline. This stability suggests that GPU memory usage of Sanitizer exhibits low sensitivity to the number of reverse engineering epochs per round. On the other hand, the time required for server-side operations per round shows a clear upward trend, escalating from 19.44 seconds at 1 epoch to 28.12 seconds at 2 epochs, and further rising to 53.73 seconds at 4 epochs. Such a trajectory in the figures stems from the inherently positive relationship between time and the number of reverse engineering epochs per round. Obviously, it persistently remains below the baseline value under varying epochs. The results show that reducing the number of round-spread reverse engineering epochs can save the server-side operation time, further validating Sanitizer’s optimization for efficiency.
@@
Impact of Different Round-spread Reverse Engineering Epochs on Effectiveness. As illustrated in Figure 9, we investigate the WDR and the ASR under different reverse engineering epochs per communication round. Specifically, both the baseline and Sanitizer consistently maintain a robust and near-perfect WDR over the set of epochs {1, 2, 4}. The ASR follows a similar trend, showing a slight and routine difference between the two approaches and remaining stable as the number of epochs varies. Results from this experiment confirm that different round-spread reverse engineering epochs have virtually no impact on the Sanitizer, which remains effective for any value of this factor. This shows that the effectiveness of Sanitizer is not sensitive to the number of round-spread reverse engineering epochs. In light of the aforementioned analysis pertaining to efficiency and effectiveness, it can be concluded that, even though Sanitizer achieves comparable effectiveness to the baseline across varying epochs, the latter is associated with significantly higher costs.
@@
Impact of Different Backdoor Subnet Sizes on Efficiency. Figure 10 reports the results. Our primary objective is to investigate the relationship of efficiency between the backdoor subnet size (across three configurations: 10%, 20%, and 40%) and the key efficiency metrics. Sanitizer exhibits an increasing trend in both GPU memory utilization and required process time as the backdoor subnet size grows. In particular, the GPU memory rises from 117 MB to 379 MB, and the process time per round extends from 16.80 seconds to 38.90 seconds. All of these results largely outperform the baseline, demonstrating the intention behind the design of Sanitizer, which aims to enhance efficiency by reducing the backdoor subnet size.
@@
Impact of Different Backdoor Subnet Sizes on Effectiveness. Similarly, we systematically vary the backdoor subnet size and investigate the impact of them on the effectiveness of Sanitizer, keeping other variables constant to ensure fairness in comparisons. Figure 10 presents the WDR results as the backdoor subnet size increases. Sanitizer achieves a near-perfect WDR of 98.77% at the smallest backdoor subnet size (10%) and maintains a high WDR across all configurations of larger subnet sizes (20% and 40%). As for the ASR, Sanitizer shows a much lower ASR than the baseline and a stable trend across three subnet sizes. Hence, while markedly reducing resource and time consumption, Sanitizer achieves comparable performance to the baseline across different backdoor subnet sizes, demonstrating its effectiveness in eliminating harmful backdoor knowledge and enabling contribution demonstration via our harmless watermarks. In summary, the effectiveness of Sanitizer is not sensitive to the reduction in subnet size within a reasonable range brought about by the need to improve efficiency during the FL process.
@@ 6 Conclusions
With the goal of strengthening trustworthy FL systems, we have introduced Sanitizer, a novel and efficient sanitization approach designed to address the harmful effects and conflict issues posed by backdoor-based client-side watermarks for contribution demonstration in practical FL applications. Sanitizer leverages backdoor subnets and round-spread reverse engineering to efficiently reverse the backdoor triggers, then confines them to their own harmless, client-specific environment before deployment and eventually makes the watermarks function in harmless ways to successfully verify the contributions. Empirical experiments have demonstrated the effectiveness, harmlessness, and especially the efficiency of Sanitizer. We advocate for further exploration in this direction and believe that Sanitizer will inspire more advanced work for better IP protection in various real-world applications.
Acknowledgement
This research is partially supported by the Croucher Start-up Allowance (Project #2499102828) and RGC Early Career Scheme (Project #27211524). Any opinions, findings, or conclusions expressed in this material are those of the authors and do not necessarily reflect the views of the Croucher Foundation and RGC.
@@ A.1 Data Availability Statements
We fully support the principles of open science, aiming to promote the transparency, reproducibility, and collaborative research. The datasets (Fashion-MNIST [74], CIFAR10 [75], CIFAR100 [75], and TinyImageNet [76]) used in this study are publicly available and widely recognized as the benchmark in the image classification tasks. The details of these datasets are as follows:
@@
Input: Total communication rounds of FL $R$; Number of clients $K$; Defense data $(x_{d},y_{d})\in\mathcal{D}_{d}$; Harmless data $x_{u}\in\mathcal{D}_{u}$; Poisoning rate of backdoor embedding $\rho\%$;
Output: Harmless watermarked FL-trained model $\mathcal{F}(\theta_{u})$.
@@
2:  for each round $r=0,1,\dots,R-1$ do
@@
9:        Conduct full Reverse Engineering on $f(\theta_{k})$ for trigger pattern $\Delta^{k}_{t}$ recovery and target class $y^{k}_{t}$ detection.        $\triangleright$ Trigger Recovery.
10:        Conduct $\min\mathbb{E}_{(\boldsymbol{x}_{d},y_{d})\in\mathcal{D}_{d}}\mathcal{L}(f\left(x_{d}+\Delta^{k}_{t};\theta_{k}\right),y_{d})$ on $f(\theta_{k})$.        $\triangleright$ Unlearning.
11:     end for
12:     Aggregation to update $\theta^{r}_{global}$ for the $r+1$ round.
13:  end for
14:  Server conducts $\min\mathbb{E}_{x_{u}\in\mathcal{D}_{u}}\mathcal{L}(f\left(x_{u}+\Delta;\theta\right),y_{t})$ on FL-trained Model $\mathcal{F}(\theta)$.       $\triangleright$ Relearning.
15:  return  Harmless watermarked FL-trained model $\mathcal{F}(\theta_{u})$.
Input: Total communication rounds of FL $R$; Number of clients $K$; Defense data $(x_{d},y_{d})\in\mathcal{D}_{d}$; Harmless data $x_{u}\in\mathcal{D}_{u}$; Rate of backdoor subnet size $s\%$; Epoch of round-spread reverse engineering per round $e$; Poisoning rate of backdoor embedding $\rho\%$;
Output: Harmless watermarked FL-trained model $\mathcal{F}(\theta_{u})$.
@@
2:  for each round $r=0,1,\dots,R-1$ do
@@
11:        $\triangleright$ Conduct round-spread reverse engineering for $e$ epochs on $f_{s}(\theta_{k}^{*})$ for trigger pattern $\Delta_{r}$ and target class $y_{r}$ at $r$ round and pruning.
@@
15:  Server conducts Harmless Relearning on FL-trained Model $\mathcal{F}(\theta)$ via Equation 7.
16:  return  Harmless watermarked FL-trained model $\mathcal{F}(\theta_{u})$.
  • Fashion-MNIST: Fashion-MNIST contains 70,000 28$\times$28 grayscale images across 10 fashion categories. The dataset is split into 60,000 training images and 10,000 test images. It is designed as a modern alternative to the classic MNIST dataset. It can be accessed from https://github.com/zalandoresearch/fashion-mnist. It is supported natively by PyTorch and will be downloaded automatically.
  • CIFAR10: CIFAR10 consists of 60,000 32$\times$32 color images in 10 different classes. The dataset is divided into 50,000 training images and 10,000 test images, and is widely used for evaluating machine learning models on small-scale object recognition tasks. CIFAR10 is publicly available at http://www.cs.toronto.edu/~kriz/cifar.html, which is supported natively by PyTorch and will be downloaded automatically.
  • CIFAR100: CIFAR100 consists of 60,000 32$\times$32 color images categorized into 100 different classes. The dataset is split into 50,000 training images and 10,000 test images, and, like CIFAR10, it is widely used for benchmarking machine learning models on small-scale object recognition tasks. It is publicly available at http://www.cs.toronto.edu/~kriz/cifar.html, and is natively supported by popular frameworks such as PyTorch, which can automatically download the dataset.
  • TinyImageNet: TinyImageNet includes 100,000 64$\times$64 color images across 200 classes. Each class contains 500 training images and 50 validation images. TinyImageNet is a smaller, more manageable subset of the full ImageNet dataset, commonly used for evaluating deep learning models. It can be downloaded from http://cs231n.stanford.edu/tiny-imagenet-200.zip.
@@
communication rounds 200
local training epochs 10
@@
Harmless learning rate 0.005
@@
weight for harmless objective 0.2
Table VIII: Default Parameter Setting in FL with Sanitizer.
@@ A.3 Default Parameter Setting
Table VIII summarizes the default parameter setting in our empirical evaluation.
@@ A.4 More Details on the Extracted Backdoor Subnet
Based on the default settings, we present a further analysis of the extracted backdoor subnet in two different architectures when applied to CIFAR10, which has a resolution of 32$\times$32 pixels and three color channels (RGB). As shown in Table IX, we compared the number of parameters and FLOPs between the original network and the extracted backdoor subnet. The results clearly demonstrate a significant reduction in computational complexity. This decrease in model size and operations highlights the efficiency gained through subnet extraction, making it more suitable for resource-constrained environments without sacrificing key performance metrics.
@@ A.5 Visual Examples of Trigger Inversion on the FL-trained Model without Sanitizer
We observe that recovering the trigger from an FL-trained model without employing Sanitizer appears to be not feasible. In this scenario, no defense mechanisms are applied during the training phase of the existing backdoor-based client-side watermarked FL, and reverse engineering is only performed on the final FL-trained model after the FL process has completed, prior to deployment. Table X illustrates that the shortcut, or trigger, associated with each class cannot be successfully recovered in the final FL-trained model. We conjecture that this outcome may be attributed to the aggregation process, which potentially dilutes the backdoor effect, thereby obscuring the true trigger. This dilution likely results from the averaging of model updates across clients, which disperses the influence of individual backdoor triggers and diminishes their impact on the global model. Consequently, the aggregated model fails to retain the malicious characteristics necessary for trigger identification. The results indicate that it seems not possible to recover the specific triggers associated with each target label, highlighting the difficulty of accurately identifying and extracting the embedded backdoors under this condition.
Table IX: Complexity of the extracted backdoor subnet in three different architectures when applied to CIFAR10.
@@
Network CIFAR10 & ViT-Base/2-32
#Params FLOPs
Original Network 16.552M 8.535G
Backdoor Subnet 639.547K 332.993M
Table X: The Original Trigger corresponding to each class and its recovered version.
@@
Table XI: Experiments with other trigger types in default settings, Blend and TrojanNN. Sanitizer also demonstrates considerable effectiveness.
Approach WDR $\uparrow$ ASR $\downarrow$ ACC $\uparrow$
Blend 95.92% 13.23% 85.46%
TrojanNN 98.63% 16.42% 82.62%
A.6 More Backdoor Trigger Types
Apart from Badnet, we conduct additional experiments to evaluate how Sanitizer performs when clients adopt different data poisoning backdoor techniques, particularly with respect to trigger types. We analyze: (i) Blend [82], which superimposes a partially transparent image as a trigger, and (ii) TrojanNN [83], which formulates an optimization problem to identify an ideal trigger pattern. Sanitizer also demonstrates considerable effectiveness. As shown in Table XI, like Badnet, Sanitizer enables watermark verification in designated, harmless environments with a success rate exceeding 95.92%, , while the same triggers can only manipulate model decisions (in a harmful manner) with a success rate of less than 16.42%. The ACCs are maintained at an appreciable level.
A.7 More Analysis and Discussion
@@
$\bullet$ Why mainly assume a Badnet-style backdoor? We argue that in a scenario where FL clients co-own the jointly trained model, clients tend to embed watermarks in the simplest and most direct manner to demonstrate their contribution, obviating the need for complex and impractical techniques. As clients cannot easily break into the source code and manipulate the standard local learning, the easiest way for them to watermark the model in FL is to conduct data poisoning backdoors because it avoids tampering with the optimization process, which many sophisticated techniques proposed for centralized learning require. This means they don’t need to expend significant effort (or resources) to demonstrate their contribution; instead, a more practical and straightforward approach for them is implementing a Badnet-like patch replacement that applies a trigger pattern onto an image and changes its label. It is reasonable to assume that they prefer to demonstrate their contribution with minimal cost rather than utilizing more sophisticated backdoor embedding methods. Moreover, we specifically focus on Badnet-like backdoor-based watermarks, as our approach is not designed to defend against backdoors in general, but rather to serve as a safeguard when such backdoors are used as watermarks under specific conditions.
A.8 Limitation and Future Work
Sanitizer has confirmed the promising results of making the backdoor-based client-side watermarks harmless to enhance the contribution demonstration in FL, but it also comes with several limitations: (a) Sanitizer has not yet optimized the computational cost associated with the subsequent step after FL, although it is relatively negligible compared to the whole FL period; (b) Sanitizer has not yet been applied to tasks beyond image classification.
In the future, we will address the challenges and further extend the applicability of Sanitizer. Sanitizer represents the first research effort that directly targets the harmlessness of backdoor-based client-side watermarks in FL. This work not only advances the state-of-the-art for image classification tasks but also paves the way for future exploration across a broader range of ML applications, including object detection and beyond. We anticipate that the principles behind Sanitizer can be effectively extended to other FL scenarios, like personalized FL, driving further innovations in enhancing the IP protection and contribution demonstration of the FL-trained model.